# EVER: Exact Volumetric Ellipsoid Rendering for Real-time View Synthesis

### v5

@@ 5.2 Changes to Adaptive Density Control
Finally, as a consequence of Section 4.2, we add an additional splitting condition: We split primitives if the opacity is near 1 across the entire primitive when viewed from the major axis to see if the gradient has vanished using the following test:
$0.99<1-\exp\mathopen{}\left(-\sigma\max\mathopen{}\left(s_{x},s_{y},s_{z}\mathclose{}\right)\mathclose{}\right)$ (8)
6 Results
Mip-NeRF360 Zip-NeRF
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
3DGS [22] 27.48 .816 .257 224 25.84 .842 .418 559
StopThePop [34] 27.33 .816 .251 180 25.92 .819 .411 403
3DGRT [30] 27.20 .818 .248 78 - - - -
SMERF [15] 27.99 .818 .238 204 27.28 .830 .389 217
Our model 27.51 .825 .233 36 26.60 .845 .368 24
ZipNeRF [4] 28.54 .828 .219 0.5 27.37 .836 .364 0.5
Table 1: Results on the Mip-NeRF360 [2] and Zip-NeRF [4] datasets. The softness of the Gaussian primitives helps with PSNR, which is a metric that prefers blurrier images because of imprecise image alignment. However, we can see from the other metrics that our method works as well as the latest 3DGS methods on the Mip-NeRF 360 datasets, and outperforms them on the Zip-NeRF datasets. Note that, unlike the prior works, our LPIPS values have been recomputed to properly normalize images to [1,1][-1,1]. Zip-NeRF outperforms all other models on all metrics, but it is prohibitively slow to render for real-time applications ($\sim\!0.5$ FPS). FPS results are computed at test resolution, and results with a were not recomputed and use different high end GPUs.
We evaluate on the 9 scenes introduced in Mip-NeRF 360 [3] and the 4 scenes introduced in Zip-NeRF [4]. We use the same parameters across all scenes, except one: the size threshold for splitting and cloning is adjusted for large scenes to avoid total densification failure for both 3DGS and our method. For the alameda scene, we utilize the exposure trick introduced in [15] to allow for evaluation on datasets with multiple exposures. We also re-tuned 3DGS to work slightly better than the tune done by Duckworth et al. [15] on the Zip-NeRF.
There have been many follow ups to 3DGS, each introducing different kinds of improvements, so we limit the scope of our comparison to best isolate our contribution. There are many recent approaches for densification [7, 41, 8, 48, 50], but we chose to change densification only in ways required to handle density-based parameterization. It is likely that recent progress on improved 3DGS heuristics may similarly improve the performance of our model. There are also many different ways to change rendering [18, 34, 5, 51, 21, 23], so we chose to only compare rendering methods that specifically address popping.
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
3DGS StopThePop Ours GT
Figure 6: Visual comparison of our method to other 3DGS-based methods on the Mip-NeRF [2] and Zip-NeRF [4] datasets. Rows 1 and 2 show regions where our results are sharper. Rows 3 and 4 show how proper blending of primitives helps with handling lighting on textureless surfaces.
6.1 Reconstruction Quality
We measure image quality through the standard image metrics PSNR, SSIM and LPIPS. Note however that we recompute LPIPS for all methods to ensure they use the same back-end (VGG) and input color range ([1,1][-1,1]). We compare our method with 3DGS [22] and Zip-NeRF [4], the current offline rendering state-of-the-art for these datasets. We also compare with StopThePop since it improves view-consistency for 3DGS, 3DGRT [30] because our work uses their ray tracing algorithm, and we SMERF [15] because even though it is a distillation based method, it is a comparable real-time 3D consistent volume rendering method.
Quantitatively, we outperform the other 3DGS-based methods across the board, and set a new state of the art in the Zip-NeRF dataset in terms of sharpness as measured by LPIPS and SSIM — even when compared to offline and distillation-based methods. Qualitatively, we notice two main factors: improved blending, and improved sharpness. The combination of more flexible primitive appearance, combined with color blending, allows our method to represent gradients better, which can be seen in Figure 4. This effect can also be seen when comparing how these methods model light falling on flat walls, as shown by the bottom two rows of Figure 6. While our method is able to reproduce the smooth image gradients in the shadowed areas, all of the Gaussian based methods struggle (similarly to how our model struggles when our blending model is ablated, as seen in Figure 8). The improved sharpness is more apparent on the larger and more difficult Zip-NeRF dataset, both in the metrics and in qualitative examples (see the second two rows of Figure 6).
Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption
Splatted No Mixing
Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption
Volume Rendering (Ours) Ground Truth
Figure 8: To visualize the difference between rendering methods, we optimize our ellipsoids for our method, then visualize how different rendering methods affect the final image. Both “Splatted” and “No Mixing” assume that Gaussians do not overlap, with the former ordering them globally for the image (like 3DGS) , and the latter sorting them independently per pixel (like StopThePop). Both approaches cause hard edges to appear in areas where different primitives were blended to produce a color gradient.
6.2 Popping
6.3 Performance
At 720p, we achieve average framerates of 36 FPS on the mip-NeRF 360 outdoor scenes (min=33), 66 FPS on the mip-NeRF 360 indoor scenes (min=57), and 30 FPS on the Zip-NeRF scenes (min=25) on an NVIDIA RTX4090. The test set resolution is high on a couple of the Zip-NeRF and Mip-NeRF360 scenes, which causes the drop in FPS in Table 1. Training takes around 1-2 hours, also on an NVIDIA RTX4090. The majority of the time is spent tracing through the scene for rendering, which takes 20-71 ms, depending on the BVH. For back propagation, 16 ms are spent storing intersections, and 13-20 ms are spent loading the primitives and atomically adding up the gradients. The BVH is rebuilt every training iteration, which takes 2-20 ms.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
@@ 6.3 Performance
Table 2: An ablation study reporting average performance on the bicycle and counter scenes from MipNeRF 360 [2] and the berlin scene from ZipNeRF [4].
@@ 6.4 Ablations
To see if it was possible to train our model using ray tracing, then render using splatting, we implemented a splatting version of our 3D ellipsoids, labeled “Splatted”. The result can be seen in Figure 8. We also tested what would happen if we sorted the primitives per ray like StopThePop , which is labeled “No Mixing”. Both of these resulted in visual degradation where the method had been mixing primitive colors together, which can be seen on the fridge and candle. These ablations also perform worse when trained for their specific rendering styles, as can be seen in Table 2. “Splatting” performs worse than “No Mixing”, which does not match the comparison between 3DGS and StopThePop, both of which have similar performance.
To ablate the effect of our densification changes, along with the other changes, we use the 3DGS renderer with the other minor changes we’ve made to densification, which we label ”3DGS+Our Changes”. These changes do not help 3DGS because they are aimed at handling density primitives. Finally, we run ablations on each of the changes we performed. ”No anisotropic” refers to ablating anisotropic loss, which results in slightly increased quality, but much lower framerate. ”No points” refers to no additional points added using the inverse contraction.”No Rand Centers” refers to ablating randomizing the pixel centers, which helps with thin structures on the bicycle scene.
@@ 7 Conclusion
We have presented Exact Volumetric Ellipsoid Rendering (EVER), which bridges the gap between slow but accurate radiance field methods such as Zip-NeRF, and fast but inaccurate radiance field methods like 3DGS. By exactly ray-tracing a volumetric collection of constant density ellipsoids, EVER is able to produce high-quality renderings that are guaranteed to be 3D-consistent and therefore exhibit no popping, and does so at 30 FPS @ 720p on a single consumer-level GPU. By combining the flexibility of ray-tracing with the speed of primitive-based radiance field methods, EVER enables highly-flexible, high-quality, real-time, radiance field reconstruction.
@@ Appendix A Backpropagation
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
3DGS StopThePop Ours GT
Figure 9: Additional visual comparison of our method to other Gaussian based methods on the Mip-NeRF 360 dataset [4].
Appendix B Full Results
PSNR berlin nyc alameda london
3DGS 26.83 26.90 24.14 25.48
StopThePop 26.81 27.14 24.12 25.61
SMERF 28.52 28.21 25.35 27.05
Our model 27.24 27.93 24.72 26.49
ZipNeRF 28.59 28.42 25.41 27.06
SSIM berlin nyc alameda london
3DGS .899 .861 .776 .830
StopThePop .885 .844 .748 .801
SMERF .887 .844 .758 .829
Our model .900 .863 .779 .837
ZipNeRF .891 .850 .767 .835
LPIPS berlin nyc alameda london
3DGS .406 .380 .441 .446
StopThePop .402 .373 .433 .438
SMERF .391 .361 .416 .390
Our model .371 .337 .389 .374
ZipNeRF .378 .331 .387 .360
@@ Appendix B Full Results
PSNR bicycle flowers garden stump treehill room counter kitchen bonsai
@@
SSIM bicycle flowers garden stump treehill room counter kitchen bonsai
@@
LPIPS bicycle flowers garden stump treehill room counter kitchen bonsai
@@
Appendix C Hyperparameters, Etc
For learning rates (LR), we change the opacity LR to 0.0125, the initial position LR to $4\times 10^{-5}$ and the final position LR to $4\times 10^{-7}$. We change the parameter known as ”percent dense” in 3DGS to $0.001785$, which controls the size threshold above which primitives are split instead of cloned. Speaking of splitting and cloning, we perform this every 200 iterations, instead of 100, and set the splitting gradient threshold to $2.5\times 10^{-7}$ and the clone gradient threshold to $0.1$. We also stop splitting and cloning at 7 million primitives, or at 16000 iterations, which ever comes first, and start at 1500 iterations.
For color, we apply softplus ($\beta=10$) to the output of the spherical harmonics, which we find avoids certain local minima where primitives get lock into a color, which happens with $\max(0,x)$ in 3DGS. Speaking of color, we increase the spherical harmonic degree every 2000 iterations, instead of 1000. One final note: we set the max primitive size to 25 units, which seems to help performance.
C.1 Inverse Contraction Initialization
To help initialize the primitives in a scene, we supplement the SfM initialization with $10000$ additional primitives. We generate these primitives by sampling their means uniformly from a radius-2 sphere. The radius is set to a constant value based the max radius at which the spheres could be packed into the radius-2 sphere, and colors are set to a constant value of 0.5. These primitives are then transformed by “uncontracting” the resulting means and covariances using the inverse of the contraction used in mip-NeRF 360 [3]. We found that highly anisotropic primitives at initialization can cause issues, so we scale the primitives to be isotropic.

### v6

@@ 6.3 Performance
Table 3: An ablation study reporting average performance on the bicycle and counter scenes from MipNeRF 360 [2] and the berlin scene from ZipNeRF [4]. The “No Anisotropic” ablation shows the cost of our regularizer that increases frame-rate.
@@ 6.4 Ablations
To see if it was possible to train our model using ray tracing, then render using splatting, we implemented a splatting version of our 3D ellipsoids, labeled “Splatted”. We also tested what would happen if we sorted the primitives per ray like StopThePop without our color blending technique, which is labeled “No Mixing”. Both of these resulted in visual degradation in areas where the method had been mixing primitive colors together, which can be seen on the fridge and candle. These ablations also perform worse when trained for their specific rendering styles, as can be seen in Table 3. “Splatting” performs worse than “No Mixing”, which does not match the comparison between 3DGS and StopThePop, both of which have similar performance.
To ablate the effect of our densification changes, along with the other changes, we use the 3DGS renderer with the other minor changes we’ve made to densification, which we label “3DGS + Our Changes”. These changes do not help 3DGS because they are aimed at handling density primitives. Finally, we run ablations on each of the changes we performed. “No anisotropic” refers to ablating anisotropic loss, which results in slightly increased quality, but much lower framerate. “No points” refers to no additional points added using the inverse contraction, which reduces the reliance on the SfM reconstruction. The inverse contraction sampling is described in the supplemental material. “No Rand Centers” refers to ablating randomizing the pixel centers, which helps with thin structures on the bicycle scene. We jitter our rays by picking a point within the pixel from a uniform random distribution, as opposed to using the center of the pixel.
@@ 7 Conclusion
Our double-intersection approach for Exact Volumetric Ellipsoid Rendering can be used to blend any number of overlapping primitive colors in a physically accurate way. We have shown that our method outperforms Zip-NeRF on the larger Zip-NeRF datasets on sharpness, while outperforming all other real-time methods on SSIM and LPIPS on the Mip-NeRF360 and Zip-NeRF datasets. By isolating the effect of our renderer, we know that this improvement comes from a combination of density based primitives and double-intersection based, physically accurate blending. This results in an approach that bridges the gap between slow but accurate radiance field methods such as Zip-NeRF, and fast but inaccurate radiance field methods like 3DGS.
We also show that sorting primitives per a pixel is not sufficient for eliminating the popping artifacts in 3DGS. It is only by combining per pixel sorting, which we achieve with ray tracing, with proper primitive blending, that we are able to eliminate these artifacts. Together, EVER is able to produce high-quality, large scale renderings that are guaranteed to be 3D-consistent, and does so at 30 FPS at 720p on a single NVIDIA RTX4090.
@@ Appendix A Backpropagation
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption      Refer to caption   Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
3DGS StopThePop SMERF Ours GT
Figure 8: Additional visual comparison of our method on the Mip-NeRF 360 dataset [4].
PSNR $\uparrow$ berlin nyc alameda london Average
3DGS 26.83 26.90 24.14 25.48 25.84
Mip Splatting 27.30 27.52 24.76 26.28 26.46
StopThePop 26.81 27.14 24.12 25.61 25.92
SMERF 28.52 28.21 25.35 27.05 27.28
Ours 27.24 27.93 24.72 26.49 26.60
Zip-NeRF 28.59 28.42 25.41 27.06 27.37
SSIM $\uparrow$ berlin nyc alameda london Average
3DGS .899 .861 .776 .830 .842
Mip Splatting .892 .853 .768 .822 .834
StopThePop .885 .844 .748 .801 .819
SMERF .887 .844 .758 .829 .830
Ours .900 .863 .779 .837 .845
Zip-NeRF .891 .850 .767 .835 .836
LPIPS $\downarrow$ berlin nyc alameda london Average
3DGS .406 .380 .441 .446 .418
Mip Splatting .392 .356 .410 .411 .392
StopThePop .402 .373 .433 .438 .411
SMERF .391 .361 .416 .390 .389
Ours .371 .337 .389 .374 .368
Zip-NeRF .378 .331 .387 .360 .364
@@
PSNR $\uparrow$ bicycle flowers garden stump treehill room counter kitchen bonsai
@@
SSIM $\uparrow$ bicycle flowers garden stump treehill room counter kitchen bonsai
@@
LPIPS $\downarrow$ bicycle flowers garden stump treehill room counter kitchen bonsai
@@
Appendix B Hyperparameters, Etc
We change the opacity learning rate to $0.0125$, the initial position learning rate to $4\times 10^{-5}$ and the final position learning rate to $4\times 10^{-7}$. We change the parameter known as “percent dense” in 3DGS to $0.001785$, which controls the size threshold above which primitives are split instead of cloned. We perform this every 200 iterations, instead of 100, and set the splitting gradient threshold to $2.5\times 10^{-7}$ and the clone gradient threshold to $0.1$. We also stop splitting and cloning at 7 million primitives, or at 16000 iterations, which ever comes first, and start at 1500 iterations.
For color, we apply a softplus activation ($\beta=10$) to the output of the spherical harmonics (instead of 3DGS’s relu activation), which we find avoids certain local minima where primitives get locked into a color. We increase the spherical harmonic degree every 2,0002,\!000 iterations, instead of 1,0001,\!000. We set the max primitive size to $25$ units, which improves performance.
B.1 Inverse Contraction Initialization
To help initialize the primitives in a scene, we supplement the SfM initialization with 10,00010,\!000 additional primitives. We generate these primitives by sampling their means uniformly from a radius-2 sphere. The radius is set to a constant value based the max radius at which the spheres could be packed into the radius-2 sphere, and colors are set to a constant value of $0.5$. These primitives are then transformed by “uncontracting” the resulting means and covariances using the inverse of the contraction used in mip-NeRF 360 [3]. We found that highly anisotropic primitives at initialization can cause issues, so we scale the primitives to be isotropic.